# SF-Loc: A Visual Mapping and Geo-Localization System based on Sparse Visual Structure Frames

### v2

@@ IV-B Map Construction and Maintenance
After the global pose optimization, we extract the frame pose from the global graph, pack the RGB image (with JPEG compression), the low-resolution depth map and the global descriptor produced by some VPR model together to construct a visual structure frame, then insert it into the map. To make the map lightweight, it is important to keep the sparsity of the visual structure frames without sacrificing too much of the effectiveness of user-side re-localization. Besides, considering the application scenario of multi-session mapping, the system should support incremental update.
Fortunately, this can be simply achieved based on the flexible model-free, frame-based map representation of SF-Loc. To be specific, the local co-visibility is checked to select new frames to be added into the map. The co-visibility is efficiently computed based on the bi-directional dense rigid flow that falls within the field of view, following
@@
where $\mathbf{u}_{ij}$ is the rigid flow in Eq. (1), $H$, $W$ are the height and width of the image. The illustration of the co-visibility is presented in Fig. 4.
The co-visibility directly reflects the overlapping of different frames. With the intention of avoiding redundant map frames, we simply select frames with relatively low co-visibility when adding frames to the map. The co-visibility threshold $\Xi$ controls the mapping process, which reflects the trade-off between data size and re-localization performance. We select $\Xi=0.4$ as the typical value, which indicates that for one query along the smooth trajectory between two nearby map frames, an optimistic $(1.0+0.4)/2=0.7$ overlapping can be expected.
When adding/removing/updating frames to the map, it is not necessary to perform data association across nearby frames or reconstruct the 3D model, as every visual structure frame works fully independently. This makes the mapping phase highly flexible, and can be easily applied to multi-session mapping or incremental map updating.
@@ V Map-Aided Localization Phase
In this phase, the implementation details of the map-aided localization are presented, which follows the coarse-to-fine process proposed in[9]. Based on state-of-the-art place recognition and feature matching techniques, specially designed methods which considers temporal integration are proposed to adapt to the map representation and to maximize the performance in real-world scenarios.
@@ V-A Multi-Frame Place Recognition
The necessity of multi-frame place recognition lies in the existence of repetitive and featureless scenes in the real world, which leads to inevitable spatial ambiguity. To handle this, it is natural to take into account the environmental cues within a larger range to resolve the ambiguity. Notice that we call this process “multi-frame place recognition” rather than “sequential place recognition”, as the map frames are generally not “sequential” but are “randomly” distributed in the geographic space.
To avoid the ambiguity of single-shot VPR, we propose an metric called spatially smoothed similarity (SSS) that combines the information of appearance similarity, map distribution and user-side relative poses. To be specific, we set particles around every landmark frames in the map, with every particle to be a 2D pose $(x,y,\theta)$ that represents a candidate of the query pose. With the help of the user-side odometry, we compute the virtual trajectory of every particle, following
@@
where $\mathcal{P}$ is the particle candidate set, $L$ is the window size used for multi-frame VPR.
For every particle, we efficiently search for the spatially nearest map frame corresponding to every history pose through KD-Tree. This leads to a sequence of L2 distances for every particle
@@
where db is the map database organized by a spatial KD-Tree, $\bm{\sigma}$ is the $L\times M$ similarity matrix which stores the L2 similarities between the $L$ query frames and all $M$ frames in the map. The similarity computation is performed every time a new query frame comes and stored in the memory.
Then, we compute the SSS distance by averaging the L2 distances
$\sigma^{i}_{\text{SSS}}=||\bm{\sigma}^{i}||/\sqrt{L},\ i\in\mathcal{P}$ (18)
and the particle with minimum $\sigma^{i}_{\text{SSS}}$ is taken as the retrieved map frame.
The rationality of SSS is based on the assumption that the historical images of a correctly recalled query should also exhibit high similarity to their spatially adjacent map frames. The geographical distribution of map frames (which is obtained in the mapping phase), along with user-side pose information, is implicitly incorporated to the metric in nearest neighbor search. This technique can be easily adopted to common VPR models without extra layers or finetuning, transforming single-frame VPR similarities to multi-frame ones.
In our implementation for outdoor environments, we simply set 3 particles around every map frame, corresponding to the 2D poses $(0,0,0^{\circ})$, $(0,0,-30^{\circ})$, $(0,0,30^{\circ})$ relative to the map frame pose.
Figure 5: Illustration of the multi-frame place recognition. Particles that indicate the query pose are set around map frames. The user-side trajectory is used for nearest searching to associate with the map frames, then the SSS distance is evaluated, which combines the information of multiple image pairs.
@@ V-B Fine Pose Estimation
After successful retrieval of the map frame corresponding to the query image, fine-grained re-localization can be performed. In our system, this process could be achieved based on one single map frame and is independent of any re-constructed model, considering that the visual structure frame contains the full-view depth. This makes the re-localization simple and highly flexible.
Specifically, 2D-2D matching between the query frame and the map frame is performed based on feature extraction method like Superpoint[53] and matching techniques like Lightglue[50] or Superglue[51]. With the 2D-2D correspondences and the full-view depth contained in the map frame, several techniques can be applied to estimate the pose of the query image, like perspective-n-point (PnP) or just aligning the two-view reconstruction result with the metric-scale depth. To improve the robustness, RANSAC can be used.
Despite the simplicity, the above methods couldn’t be directly used for multi-frame matching. In the application scenarios that our system concerns (cross-season, dynamic objects, etc.), robust feature matching is hard to promise, even with state-of-the-art learning-based techniques. To maximize the success rate and provide accurate pose estimation under large amounts of outliers, we combine multi-frame correspondences and the odometry in a factor graph, as shown in Fig. 6. For better efficiency, we don’t do exhaustive matching among all the related frames. Instead, we compute the correspondences between every query-to-database image pair from VPR. Based on the re-projection errors, the cost function of the multi-frame pose estimation is written as follows
$\begin{split}\min\!\sum_{k\in\mathcal{K}}\!\!\left\|{\mathbf{T}^{w}_{c_{k+1}}}^{\!\!\!\!\!\!-1}\!\!\circ\!\mathbf{T}^{w}_{c_{k}}\!\!\circ\!\hat{\mathbf{T}}^{c_{k}}_{c_{k+1}}\!\right\|^{2}_{\bm{\Omega}_{p}}\!\!\!\!+\!\!\!\sum_{k\in\mathcal{K}}\!\sum_{f\in\mathcal{F}_{k,i}}\!\!\!\!\left\|\lambda^{m_{i}}_{f}\!\!-\!\!\hat{\lambda}^{m_{i}}_{f}\!\right\|^{2}_{\Omega_{\lambda}}\\ \!\!\!+\!\!\sum_{k\in\mathcal{K}}\!\sum_{f\in\mathcal{F}_{k,i}}\!\!\!\rho_{C}\!\!\left(\left\|\pi_{c}\!\left(\!\mathbf{T}^{c_{k}}_{m_{i}}\!\!\circ\!{\pi}_{m}^{-1}(\mathbf{u}^{m_{i}}_{f},\lambda^{m_{i}}_{f})\right)\!-\!\mathbf{u}^{c_{k}}_{f}\right\|^{2}_{\bm{\Omega}_{c}}\right)\end{split}$ (19)
where $\hat{\mathbf{T}}^{c_{k}}_{c_{k\!+\!1}}$ is the relative pose estimation, $\hat{\lambda}^{m_{i}}_{f}$ is the sampled inverse depth of point $f$ on the map frame $m_{i}$, $\mathcal{F}_{k,i}$ is the matched feature point set between map frame $m_{i}$ and user-side frame $c_{k}$, $\mathbf{u}^{m_{i}}_{f}$ and $\mathbf{u}^{c_{k}}_{f}$ are the point coordinates respectively, ${\bm{\Omega}_{p}}$, ${\Omega_{\lambda}}$, ${\bm{\Omega}_{c}}$ are the uncertainties. Notice that the map frame poses are fixed due to unobservability of the reference.
@@
Another concern about the re-localization is that, does the depth estimated by monocular DBA enough for high-precision re-localization? A rough answer is that the fine-grained visual localization doesn’t need decimeter-level depth to achieve decimeter-level pose estimation. An intuitive example is illustrated in Fig. 7. As is shown, the re-localization error is proportional to the product of the depth (or rather scale) error and the recall distance. Given 10% depth error of the map frame and $<$ 10 m recall distance, we can still expect a $<$ 1 m re-localization accuracy, which is enough for the applications that we care about. Actually, considering the demand of robust re-localization, correct image retrieval and feature matching are more important than the accuracy of structure information, which allow us to store the depth of the map frame in a highly compact way (i.e., $\times$8 downsampled).
In the later experimental part, we will discuss the performance of different pose estimation methods.
@@ VI Experiments
Figure 8: Vehicle trajectories and example images of the experimental dataset. The data sequence collected in 2022/10/23 (5400 sec duration, 30 km mileage, 16 km one-way road coverage) is used for mapping, while the data sequence collected in 2023/04/12 (1800 sec duration, 10 km mileage, 4.9 km one-way road coverage) is used for localization.
In the experimental part, different modules of the system, including multi-sensor mapping, place recognition and fine re-localization are comprehensively tested.
To fully evaluate the mapping and localization performance of the system in complex conditions, we use the self-made multi-sensor dataset collected in Wuhan City on 2022/10/23 and 2023/04/12 respectively. The trajectories of the two sequences are shown in Fig. 8. To be specific, the data sequence on 2022/10/23 is used to evaluate the performance of incremental mapping, with multiple revisits of the same road segments. The data sequence on 2023/04/12 is used to evaluate the map-aided localization, which is challenging considering the cross-season condition and the high vegetation coverage rate of the environment.
The experimental vehicle is equipped with RGB cameras, an ADIS16470 IMU, a Septentrio AsteRx4 GNSS receiver and a tactical-grade IMU for reference use. The sensor installation and the camera lens are different for the two dates. A nearby base station is used for differential GNSS (DGNSS) processing. The smoothed trajectory of post-processing DGNSS/INS integration based on high-end IMU is taken as the ground truth, which is featured with centimeter-to-decimeter level position accuracy.
In the mapping phase, 5 Hz images (512 $\times$ 384 ) are used. In the localization phase, 1 Hz images (512 $\times$ 384 ) are used; For multi-frame localization, except for the current frame, almost stationary frames are skipped.
@@ VI-A Mapping Performance
Figure 9: Pose estimation results of the mapping phase. Monocular schemes are used for vision-based methods. The bold numbers are root mean square errors (RMSE) of the horizontal position error.
We first focus on the effectiveness of the mapping phase.
For large-scale mapping, obtaining precise global poses is fundamental. Nevertheless, this is a challenging task considering the intermittent degradation of GNSS observations in the experimental scenario. In Fig. 9, we show the pose estimation results of different approaches, including: 1) GNSS RTK, 2) VINS-Fusion (monocular VIO + RTK, final optimized trajectory) and 3) the proposed multi-sensor DBA scheme in real-time mode and global mode respectively. As can be found, the GNSS RTK results are noisy. Although decimeter-level accuracy is intermittently met, gross errors occur frequently. By integrating RTK with VIO, the VINS-Fusion scheme suppresses the gross errors to some extent. For the proposed DBA (real-time) scheme, by integrating visual DBA, inertial information and RTK in a tightly-coupled way, the GNSS errors are better mitigated. Furthermore, with the global optimization presented in Sect. IV-A, the position estimation error the DBA (global) scheme is mostly kept to below 1 m throughout the sequence, which ensures the decimeter-level lower bound of user-side absolute localization.
TABLE I: Map storage of Different Mapping Schemes. In Calculating the Distance, Non-Repeating One-Way Roads are Considered.
Map Scheme Storage (MB) Storage per Km (MB) Frame Number Average Distance (m)
ORBSLAM3 6590.51  1 403.53 18768 0.870
HLoc (SP) 32206.7  2 1971.99 27026 0.607
SF-Loc (0.4) 48.840  3 2.990 1743 9.37
SF-Loc (0.3) 33.438  4 2.047 1189 13.74
@@
    3 37.67 MB (JPEG, 60) + 3.41 MB (Global Desc.) + 7.76 MB (Depth)
    4 25.81 MB (JPEG, 60) + 2.32 MB (Global Desc.) + 5.30 MB (Depth)
@@
Figure 10: Example cases of the coarse localization. The green borderline indicates correct retrieval, while the red borderline indicates wrong retrieval (large position error or heading error). High-similarity candidates (local minima and $\sigma<{\sigma}_{\text{min}}+0.15({\sigma}_{\text{max}}-{\sigma}_{\text{min}})$) are marked using larger circles.
@@
On the other hand, the generation of map data is looked into. As baselines, we test ORBSLAM3 and HLoc for the mapping task, which are both commonly used pipelines for visual mapping and re-localization. For ORBSLAM3, we simply use the monocular visual-inertial mode to process the sequence and save the map data in the Atlas format. The map storage is over 6 GB for the whole sequence, which is relatively heavy for practical applications. As for HLoc, we found it hard to achieve such large-scale SfM through the default COLMAP[54] + SIFT[55] pipeline. Instead, we directly use the ground-truth camera poses and perform feature extraction, matching and triangulation based on Superpoint and SuperGlue. The map storage is over 30 GB, which is 5 times larger than the ORBSLAM3 map. Such heavy map makes it hard to be applied to large-scale scenarios.
For the proposed SF-Loc, after global optimization, the visual structure frames are sparsified through co-visibility checking and inserted into the global map. Under such mechanism, the map is only updated when new viewpoints are available, and the number of the map frames wouldn’t grow if the vehicle revisited the region several times. As mentioned in Sect. IV-B, the co-visibility threshold $\Xi$ directly controls the map sparsity. We test different co-visibility thresholds (0.3 or 0.4) for map construction. Finally, the data size of the map is around 3 MB per kilometer for $\Xi=0.4$ and around 2 MB per kilometer for $\Xi=0.3$, which is 1$\sim$3 orders of magnitude smaller than that of the aforementioned methods. Later, we will discuss how high-precision localization can be achieved under such sparsified and compressed map information.
@@
TABLE II: Recall and Accuracy of the Coarse Re-Localization. Map Frames with Different Sparsities ($\Xi=0.4$ and $\Xi=0.3$) are Considered Respectively.
$\Xi$ Method Recall @5 m Recall @10 m Recall @20 m RMSE1 (m)
0.4 ORB + DBoW2 1.83% 3.11% 4.39% 38.79
NetVLAD 43.09% 59.36% 65.35% 15.51
AnyLoc 56.98% 79.09% 89.52% 13.91
MixVPR 65.91% 89.67% 96.39% 7.67
CosPlace18 64.35% 89.17% 94.67% 7.87
CosPlace50 69.57% 93.34% 97.45% 7.12
EigenPlaces18 69.07% 91.45% 96.17% 7.04
EigenPlaces50 70.74% 92.73% 96.95% 6.52
EigenPlaces18 (5) 69.35% 91.95% 96.67% 7.00
EigenPlaces50 (5) 70.91% 93.28% 97.56% 6.54
EigenPlaces18 (10) 69.52% 92.23% 96.95% 6.93
EigenPlaces50 (10) 71.90% 94.50% 98.83% 5.96
EigenPlaces18 (SSS5) 72.68% 96.72% 99.83% 5.18
EigenPlaces50 (SSS5) 75.62% 97.06% 99.89% 4.58
EigenPlaces18 (SSS10) 73.96% 97.67% 100.00% 4.55
EigenPlaces50 (SSS10) 73.57% 97.89% 100.00% 4.52
0.3 ORB + DBoW2 1.17% 2.28% 3.11% 45.11
NetVLAD 34.37% 53.41% 61.41% 16.40
AnyLoc 48.45% 75.99% 87.78% 13.92
MixVPR 53.91% 85.34% 94.84% 9.07
CosPlace18 52.47% 82.51% 92.67% 8.74
CosPlace50 56.08% 89.28% 96.50% 8.32
EigenPlaces18 54.36% 85.12% 93.73% 8.05
EigenPlaces50 56.75% 89.84% 97.06% 7.50
EigenPlaces18 (5) 54.75% 85.79% 94.50% 7.95
EigenPlaces50 (5) 57.13% 90.28% 97.61% 8.75
EigenPlaces18 (10) 54.91% 86.12% 94.95% 7.79
EigenPlaces50 (10) 57.25% 90.56% 97.95% 7.00
EigenPlaces18 (SSS5) 59.19% 92.84% 99.44% 5.73
EigenPlaces50 (SSS5) 59.74% 93.56% 99.83% 5.63
EigenPlaces18 (SSS10) 57.97% 92.84% 99.22% 5.50
EigenPlaces50 (SSS10) 58.80% 93.78% 100.00% 5.48
  • 1 For RMSE calculation, only $<$ 100 m cases are taken into account.

### v3

@@ IV-B Map Construction and Maintenance
After global pose optimization, we extract the frame pose from the global graph and form a visual structure frame by combining the JPEG-compressed RGB image, low-resolution depth map, and global descriptor from the VPR model, which is subsequently inserted into the map. To make the map lightweight, it is important to keep the sparsity of the visual structure frames without significantly compromising re-localization performance. Besides, considering the application scenario of multi-session mapping, the map should support flexible extension and update.
Fortunately, this can be simply achieved based on the model-free, frame-based map representation of SF-Loc. To be specific, the local co-visibility is checked to preserve essential frames in the map. With the global poses and dense depths, the co-visibility is efficiently computed based on the bi-directional dense rigid flow that falls within the field of view, following
@@
where $\mathbf{u}_{ij}$ is the computed rigid flow in Eq. (1), $H$, $W$ are the height and width of the image. The illustration of the co-visibility is presented in Fig. 4.
The co-visibility directly reflects the overlapping of different frames. With the intention of avoiding redundant map frames, the inter-frame co-visibility is controlled under a threshold $\Xi$. Considering adding/removing/updating frames, the following mechanism is employed:
1) When the mapping vehicle passes through a place and generates a new map frame, existing map frames within a certain range nearby are retrieved.
2) The bidirectional co-visibility is checked between the new map frame and other frames.
3) If the bidirectional co-visibility between the new map frame and all other frames is below $\Xi$, add the new map frame to the map.
4) If there exist map frames with bidirectional co-visibility above $\Xi$ with the new map frame, and the score of the new map frame is higher than that of the existing frames, delete the existing frames and add the new map frame to the map, which actually update the map locally; otherwise, discard the new map frame.
Here, the score can refer to metrics such as freshness, uncertainty or the proportion of dynamic objects.
It is noted that the map update mechanism is not designed to achieve the most complete scene coverage but to provide relatively large coverage with low redundancy. Under different considerations, different score metrics and update policies can be employed. The co-visibility threshold $\Xi$ directly controls the map sparsity, which reflects the trade-off between data size and re-localization performance. We select $\Xi=0.4$ as the typical value, which indicates that for one query along the smooth trajectory between two nearby map frames, an optimistic $(1.0+0.4)/2=0.7$ overlapping can be expected.
When adding/removing/updating frames in the map, it is not necessary to perform data association across nearby frames or reconstruct the 3D model, as every visual structure frame works fully independently. This makes the mapping phase highly flexible, and can be easily applied to multi-session mapping or incremental map updating.
@@ V Map-Aided Localization Phase
In this section, the implementation details of the map-aided localization are presented, which follows the coarse-to-fine process like [9]. Based on state-of-the-art place recognition and feature matching models, specially designed methods which consider spatiotemporal association are proposed to align with the map representation and to maximize the performance in realistic scenarios.
@@ V-A Multi-Frame Place Recognition
The necessity of multi-frame place recognition lies in the existence of repetitive and featureless scenes in the real world, which leads to inevitable spatial ambiguity. To handle this, it is natural to take into account the environmental cues within a larger range to resolve the ambiguity. Notice that we call this process “multi-frame place recognition” rather than “sequential place recognition”, as the map frames are generally not “sequential” but are “randomly” distributed in the geographic space. Especially for multi-session mapping or at intersections, the spatially nearby map frames may come from multiple non-consecutive sequences.
To avoid the ambiguity of single-shot VPR, we propose a metric called spatiotemporally associated similarity (SAS) that combines the information of appearance similarity, map distribution and user-side relative poses. To be specific, we set particles around every landmark frames in the map, with every particle to be a 2D pose $(x,y,\theta)$ that represents a candidate of the query pose. With the help of the user-side odometry, we compute the virtual trajectory of every particle, following
@@
where $\mathbf{T}^{w}_{p_{i}}$ is the particle pose, $\left(\mathbf{T}^{i}\right)_{4\times 4\times L}$ is the virtual trajectory, $\mathcal{P}$ is the set of particle candidates, $L$ is the window size used for multi-frame VPR.
For every virtual trajectory, we efficiently search for the spatially nearest map frame corresponding to every history pose through KD-Tree, as shown in Fig. 5. This leads to a sequence of L2 distances for every particle
@@
where db is the map database organized by a spatial KD-Tree for fast searching, $\bm{\sigma}$ is the $L\times M$ similarity matrix which stores the L2 similarities between the $L$ query frames and all $M$ frames in the map, $\cdot$]” is the indexing operator. The similarity computation is performed every time a new query frame comes and stored in the memory,
Then, we compute the SAS distance by averaging the L2 distances
[…]
and the particle with minimum $\sigma^{i}_{\text{SAS}}$ is taken as the retrieved map frame.
The rationale of SAS is based on the assumption that the historical images of a correctly relocalized query should also exhibit high similarity to their spatially adjacent map frames, which helps overcome the multipeaked characteristic of descriptor-based place similarity evaluation function. The geographical distribution of map frames, along with user-side pose information, is implicitly incorporated into the metric in nearest neighbor search. In this way, the spatiotemporal association is naturally achieved, without the need of an exact sequence-to-sequence matching. This technique can be easily adopted to common VPR models without extra layers or finetuning, transforming single-frame VPR similarities to multi-frame ones, which effectively broaden the perceived context for more discriminative recognition.
In our implementation for outdoor environments, we simply set 3 particles around every map frame, corresponding to the 2D poses $(0,0,0^{\circ})$, $(0,0,-30^{\circ})$, $(0,0,30^{\circ})$ relative to the map frame pose, which aligns with the 60 camera FOV and helps overcome the possible heading inconsistency of user/map frame poses.
Figure 5: Illustration of the multi-frame place recognition. Particles that indicate the query pose are set around map frames. The user-side trajectory is used for nearest searching to associate with the map frames, then the SAS distance is evaluated, which combines the information of multiple image pairs.
@@ V-B Fine Pose Estimation
After successful retrieval of the map frame corresponding to the query image, fine-grained re-localization can be performed.
Specifically, 2D-2D matching between the query and map frames is performed based on either feature-based methods[43] or detector-free methods[48]. With these correspondences and the full-view depth contained in the map frame, several techniques can be applied to estimate the pose of the query image, like perspective-n-point (PnP) or just aligning the two-view reconstruction result with the metric-scale depth.
Despite the simplicity, the above methods couldn’t be directly used for multi-frame matching. In the application scenarios that our system concerns (cross-temporal, significant appearance variation, dynamic objects, etc.), robust feature matching is hard to promise. To provide accurate pose estimation under numerous outliers, we combine multi-frame correspondences and the odometry in a factor graph, as shown in Fig. 6. For better efficiency, we don’t do exhaustive matching among all the related frames. Instead, we compute the correspondences between every query-to-database image pair from VPR. Based on the re-projection errors, the cost function of the multi-frame pose estimation is written as follows
$\begin{split}\min\!\sum_{k\in\mathcal{K}}\!\!\left\|{\mathbf{T}^{w}_{c_{k+1}}}^{\!\!\!\!\!\!-1}\!\!\circ\!\mathbf{T}^{w}_{c_{k}}\!\!\circ\!\tilde{\mathbf{T}}^{c_{k}}_{c_{k+1}}\!\right\|^{2}_{\bm{\Omega}_{p}}\!\!\!\!+\!\!\!\sum_{k\in\mathcal{K}}\!\sum_{f\in\mathcal{F}_{k,i}}\!\!\!\!\left\|\lambda^{m_{i}}_{f}\!\!-\!\!\tilde{\lambda}^{m_{i}}_{f}\!\right\|^{2}_{\Omega_{\lambda}}\\ \!\!\!+\!\!\sum_{k\in\mathcal{K}}\!\sum_{f\in\mathcal{F}_{k,i}}\!\!\!\rho_{C}\!\!\left(\left\|\Pi_{c}\!\left(\!\mathbf{T}^{c_{k}}_{m_{i}}\!\!\circ\!{\Pi}_{m}^{-1}(\mathbf{u}^{m_{i}}_{f},\lambda^{m_{i}}_{f})\right)\!-\!\mathbf{u}^{c_{k}}_{f}\right\|^{2}_{\bm{\Omega}_{c}}\right)\end{split}$ (19)
where $\tilde{\mathbf{T}}^{c_{k}}_{c_{k\!+\!1}}$ is the relative pose estimation from the odometry, $\tilde{\lambda}^{m_{i}}_{f}$ is the sampled inverse depth of point $f$ on the fixed map frame $m_{i}$, $\mathcal{F}_{k,i}$ is the matched feature point set between map frame $m_{i}$ and user-side frame $c_{k}$, $\mathbf{u}^{m_{i}}_{f}$ and $\mathbf{u}^{c_{k}}_{f}$ are the point coordinates respectively, $\bm{\Omega}_{p}$}, $\Omega_{\lambda}$}, $\bm{\Omega}_{c}$} are the uncertainties.
@@
Another concern about the re-localization is whether the depth derived by monocular DBA sufficient for precise re-localization. An intuitive example is illustrated in Fig. 7. As is shown, the re-localization error is proportional to the product of the depth error (simplified as scale error) and the recall distance. In vehicle-based multi-view triangulation, it is a moderate assumption to expect <<10% depth estimation uncertainty for objects within 50 m distance in most regions of the field of view (FOV). Given 10% depth error of the map frame and <<10 m recall distance, we can still expect a <<1 m re-localization accuracy, which is adequate for our purposes. For the demand of robust re-localization, correct image retrieval and feature matching matter more than the accuracy of structure information, allowing us to store the map frame depth in a highly compact way (i.e., $\times$8 downsampled, which naturally aligns with the DBA process Eq. (1)).
@@ VI Experiments
Figure 8: Vehicle trajectories and example images of the experimental dataset. Three 1800-sec data sequences collected in 2022/10/23 (30 km mileage, 16 km one-way road coverage in total) are used for multi-session mapping, while the 1800-sec data sequence collected in 2023/04/12 (10 km mileage, 4.9 km one-way road coverage) is used for localization. Different trajectories are offset in the Z-direction to provide clearer visualization.
In the experimental part, different modules of the system, including multi-sensor mapping, place recognition and fine re-localization are comprehensively tested.  [^2] All pretrained models and datasets are used under their respective licenses.
To fully evaluate the mapping and localization performance of the system in complex conditions, we use the self-made dataset collected in Wuhan City on 2022/10/23 and 2023/04/12, featuring large scale, complete sensor coverage and reliable ground truth. Specifically, three data sequences on 2022/10/23 are used to evaluate the performance of incremental mapping, which provides relatively complete coverage of the region’s road segments and exhibits multiple revisits. The data sequence on 2023/04/12 is used to evaluate the map-aided localization. Some challenging conditions are shown in Fig. 8, including exposure difference, occlusion, scene change and repetitive patterns.
The experimental vehicle is equipped with RGB cameras, an ADIS16470 IMU, a Septentrio AsteRx4 GNSS receiver and a tactical-grade IMU for reference use. The FOVs of the images used for mapping and localization are different (60 vs. 57 horizontally). A nearby base station is used for differential GNSS (DGNSS) processing. The smoothed trajectory of post-processing DGNSS/INS integration based on high-end IMU is taken as the ground truth, which is featured with centimeter-to-decimeter level position accuracy.
In the mapping phase, 5 Hz images (512 $\times$ 384) are used. In the localization phase, 1 Hz images (512 $\times$ 384) are used. For multi-frame localization, except for the current frame, almost stationary frames are skipped. Considering the practical requirements of geo-localization applications, the focus is solely on horizontal positioning accuracy.
@@ VI-A Mapping Performance
Figure 9: Pose estimation results of the mapping phase. Monocular schemes are used for vision-based methods. The bold numbers are root mean square errors (RMSE) of the horizontal position estimation.
Figure 10: Multi-session mapping results (co-visbility threshold $\Xi=0.4$) with different policies. (a) Incremental: In this mode, the frames of newer sessions have same scores with the old frames, thus the map would only be extended when new viewpoints are available. (b) Freshness first: In this mode, the frames of newer sessions have higher scores than the old frames, and the map would be updated when the vehicle passes through pre-built roads in newer sessions.
We first focus on the effectiveness of the mapping phase. The data of different sessions are processed separately.
For large-scale mapping, obtaining accurate global poses is fundamental. Nevertheless, this is a challenging task considering the intermittent degradation of GNSS observations in the experimental scenario. In Fig. 9, we show the pose estimation results of different approaches, including: 1) GNSS RTK, 2) VINS-Fusion (monocular VIO + RTK, final optimized trajectory) and 3) the proposed multi-sensor DBA scheme in real-time mode and global mode respectively.
The results shown in Fig. 9 demonstrate that the DBA (global) scheme offers effective resilience to GNSS and VIO degradation, providing improved performance compared to VINS-Fusion and DBA (real-time) with RMSE reductions of 75.0% and 45.2% respectively. With the global optimization presented in Sect. IV-A, the position estimation error the DBA (global) scheme is mostly kept to decimeter-level throughout the sequence, which ensures the reference accuracy for user-side absolute localization.
Based on the global poses and dense depths, co-visibility-based multi-session mapping is performed. The distribution of generated map frames following different policies are shown in Fig.10. The results demonstrate the effectiveness and flexibility of the proposed method in performing low-redundancy mapping under different criteria. In later discussions, we simply use the map generated in “incremental” mode for evaluation.
TABLE I: Map storage of Different Mapping Schemes. For SF-Loc, Different Co-Visibility Thresholds ($\Xi=0.4/0.3$) are Considered. In Calculating the Distance, Non-Repeating One-Way Roads are Considered.
Map Scheme
Storage
Storage
per Km (MB)
Frame
Number
Average
Distance (m)
ORBSLAM3
6590.51
1
403.53
18768
0.870
HLoc (SuperPoint)
32206.7
2
1971.99
27026
0.607
SF-Loc ($\Xi$=0.4)
48.473
3
2.967
1727
9.46
SF-Loc ($\Xi$=0.3)
33.007
4
2.021
1172
13.94
@@
    3 37.37 MB (JPEG, 60) + 3.37 MB (Global Desc.) + 7.72 MB (Depth)
    4 25.47 MB (JPEG, 60) + 2.29 MB (Global Desc.) + 5.24 MB (Depth)
@@
Figure 11: Example cases of the coarse localization ($\Xi=0.4$). The green borderline indicates correct retrieval, while the red borderline indicates wrong retrieval (large position error or heading error). High-similarity candidates (local minima and $\sigma<{\sigma}_{\text{min}}+0.15({\sigma}_{\text{max}}-{\sigma}_{\text{min}})$) are marked using larger circles.
@@
Furthermore, the generated map data is analyzed. As baselines, we test ORBSLAM3 and HLoc for comparison, which are both commonly used pipelines for visual mapping and re-localization. For methods based on neural/learnable representations (e.g., NeRF [11], 3DGS[37, 12]), as existing implementations lack support for such large-scale data (30 km length) and are not directly applicable to geo-localization tasks, they are not included in the evaluation. For ORBSLAM3, we use the monocular visual-inertial mode to process the sequence and save the map data in the Atlas format. The map storage is over 6 GB for the whole sequence, which is relatively heavy for practical applications. As for HLoc, we found it hard to achieve such large-scale SfM through the default COLMAP[46] pipeline. Instead, we directly use the ground-truth camera poses and perform feature extraction, matching and triangulation based on Superpoint[42] and SuperGlue[44]. The map storage is over 30 GB, which is 5 times larger than the ORBSLAM3 map. Such heavy map makes it hard to be applied to large-scale scenarios.
[…]
@@
TABLE II: Recall and Accuracy of the Coarse Re-Localization. Maps with Different Sparsities (Co-Visbility Threshold $\Xi=0.4/0.3$) are Considered Respectively.
$\Xi$
Method
Recall
@5 m
Recall
@10 m
Recall
@20 m
RMSE1
0.4 ORB + DBoW2
1.28%
2.01%
2.62%
49.17
NetVLAD
45.19%
59.53%
67.46%
15.01
AnyLoc
59.10%
80.39%
89.73%
12.99
MixVPR
65.89%
86.75%
95.60%
7.58
CosPlace18
59.86%
84.41%
95.38%
8.22
CosPlace50
61.54%
85.71%
95.55%
8.57
EigenPlaces18
61.43%
84.57%
94.62%
8.35
EigenPlaces50
63.88%
87.62%
97.12%
7.77
EigenPlaces18 (5)
61.71%
85.44%
96.31%
9.28
EigenPlaces50 (5)
64.31%
88.32%
98.21%
7.18
EigenPlaces18 (10)
61.76%
85.61%
96.47%
8.75
EigenPlaces50 (10)
64.48%
88.76%
98.75%
7.00
EigenPlaces18 (SAS5)
74.69%
95.76%
99.57%
6.07
EigenPlaces50 (SAS5)
74.74%
96.36%
99.89%
4.78
EigenPlaces18 (SAS10)
74.47%
97.01%
100.00%
4.65
EigenPlaces50 (SAS10)
74.58%
96.80%
100.00%
4.57
0.3 ORB + DBoW2
1.16%
1.77%
2.80%
47.78
NetVLAD
34.11%
51.44%
61.60%
17.45
AnyLoc
47.26%
76.10%
88.21%
13.82
MixVPR
52.69%
84.46%
95.27%
8.32
CosPlace18
50.03%
82.24%
94.79%
8.88
CosPlace50
50.30%
82.94%
95.17%
9.57
EigenPlaces18
51.66%
83.76%
94.73%
8.26
EigenPlaces50
53.50%
86.04%
96.80%
8.25
EigenPlaces18 (5)
51.93%
84.41%
95.60%
9.49
EigenPlaces50 (5)
53.72%
86.91%
97.94%
7.97
EigenPlaces18 (10)
51.71%
84.36%
95.71%
8.44
EigenPlaces50 (10)
53.56%
86.64%
97.61%
7.96
EigenPlaces18 (SAS5)
55.78%
91.80%
99.02%
6.70
EigenPlaces50 (SAS5)
56.60%
92.29%
99.57%
6.01
EigenPlaces18 (SAS10)
56.93%
92.50%
99.78%
5.97
EigenPlaces50 (SAS10)
56.49%
92.45%
100.00%
5.80
  • 1 For RMSE calculation, only << 100 m cases are taken into account.
TABLE III: Recall of the Coarse Re-Localization with Different Multi-Frame Processing Techniques based on Different VPR Models. The SF-Loc Map with $\Xi=0.4$ is Considered.
VPR
Model
Multi-
Frame
Recall
@5 m (%)
Recall
@10 m (%)
Recall
@20 m (%)
NetVLAD -
45.19
59.53
67.46
10
49.05 (3.86↑)
66.11 (6.58↑)
76.75 (9.29↑)
SAS10
69.20 (24.0↑)
91.31 (31.8↑)
96.31 (28.9↑)
AnyLoc -
59.10
80.39
89.73
10
60.24 (1.14↑)
82.07 (1.48↑)
92.02 (2.29↑)
SAS10
71.92 (12.8↑)
95.60 (15.2↑)
99.51 (9.78↑)
MixVPR -
65.89
86.75
95.60
10
66.16 (0.27↑)
87.29 (0.54↑)
96.52 (0.92↑)
SAS10
72.30 (6.41↑)
96.36 (9.61↑)
99.95 (4.35↑)